# Riemannian Flow Matching for Brain Connectivity Matrices via Pullback Geometry

### v1

@@ 4.3 Baselines
These baselines apply generative models to the lower-triangular part of SPD or correlation matrices, a common heuristic in manifold modeling [44]. For fMRI, they use the strictly lower-triangular entries ($\phi:{\text{Corr}_{d}}\mapsto\mathbb{R}^{d(d-1)/2}$); for EEG, all lower-triangular entries ($\phi:{\mathbb{S}_{d}^{++}}\mapsto\mathbb{R}^{d(d+1)/2}$). TriangDDPM uses a DDPM [2], while TriangCFM trains a vector field using the standard CFM loss. Since $\phi$ is not diffeomorphic, generated vectors that do not belong to the manifold are projected back onto it (see Appendix L). These methods trade geometric fidelity for simplicity, relying on post hoc projections to enforce constraints.
@@
We apply the log–Euclidean map $\phi_{\mathbb{S}_{d}^{++}}$ (9) for EEG and the normalized Cholesky map $\phi_{\text{Corr}_{d}}$ (12) for fMRI in Algorithm 1 and 2. These diffeomorphisms allow Euclidean training with CFM while ensuring manifold-valid samples without post-processing. Implementation details are presented in Appendix J.
@@ 5 Results
We report both quantitative results and a neurophysiological plausibility study. Quantitative comparisons are summarized in Table 1, with additional analysis of projection effects in Table 2. Neurophysiological relevance is assessed in Figure 2, which shows class-conditional fMRI connectomes via Fréchet means and topographic maps of EEG Common Spatial Patterns (CSP) filters. Together, these results show that DiffeoCFM produces realistic, class-conditional samples that preserve key features of brain connectivity. For a more detailed discussion of the baselines, please refer to Appendix G.
@@ 5.1 Quantitative Study
As shown in Table 1, DiffeoCFM consistently outperforms all baseline generative models across datasets in terms of $\alpha$-precision and $\alpha,\beta$-F1, establishing itself as the most robust method for aligning with the true data distribution. It also achieves the highest $\beta$-recall across the three fMRI datasets. On EEG datasets, DiffeoGauss achieves slightly higher $\beta$-recall but at the cost of much lower $\alpha$-precision, leading to weaker overall $\alpha,\beta$-F1 scores. In contrast, TriangDDPM and TriangCFM perform poorly on all quality metrics after projection, due to a substantial degradation in sample quality. As detailed in Table 2, projecting onto the manifold—${\text{Corr}_{d}}$ for fMRI and ${\mathbb{S}_{d}^{++}}$ for EEG—introduces a significant performance drop. This is because TriangDDPM and TriangCFM generate matrices that visually resemble realistic connectivity patterns but contain negative eigenvalues, making them invalid. The projection step corrects these matrices but distorts their structure, leading to sharp decreases in $\alpha,\beta$-F1—up to $-0.74$ on ADNI and $-0.76$ on OASIS-3—rendering TriangCFM impractical for use. This degradation is also visible in Figure 4, where post-projection alter the structure of fMRI connectomes (see Appendix L). Finally, on EEG datasets, DiffeoCFM nearly matches the performance of Real Data in terms of $\alpha,\beta$-F1, suggesting high sample realism.
Table 2: Impact of projection onto the manifolds ${\mathbb{S}_{d}^{++}}$ and ${\text{Corr}_{d}}$: performance difference Δ=TriangCFMTriangCFM without $\Delta=\text{TriangCFM}-\text{TriangCFM without projections}$. Negative $\Delta$ indicate degraded sample quality after projection onto the manifold. Projection enforces geometric constraints but severely degrades sample fidelity.
Dataset $\alpha$-precision $\beta$-recall $\alpha$,$\beta$-F1 ABIDE -0.34 -0.69 -0.50 ADNI -0.63 -0.74 -0.69 OASIS-3 -0.52 -0.76 -0.64 BNCI 2014-002 +0.13 -0.56 -0.19 BNCI 2015-001 +0.00 -0.19 -0.09
@@
For the CAS metrics, which evaluate downstream predictive performance using ROC-AUC and F1 scores, DiffeoCFM consistently achieves strong results, often approaching the performance of Real Data. It obtains the highest ROC-AUC and F1 scores across all fMRI and EEG datasets. While DiffeoGauss remains competitive in terms of ROC-AUC, DiffeoCFM substantially outperforms it on F1 scores, with absolute gains of +0.08, +0.19, and +0.13 on ABIDE, ADNI, and OASIS-3, respectively. TriangDDPM and TriangCFM perform poorly across both CAS metrics on all datasets. This underperformance reflects the effect of the projection step, which alters generated matrices in ways detrimental to downstream classification. It is important to note that the classification pipeline requires inputs to be valid SPD matrices (i.e., elements of ${\mathbb{S}_{d}^{++}}$). As a result, CAS metrics cannot be computed for TriangDDPM and TriangCFM without projection. For this reason, their performance without projection is not reported in Table 2.
@@
To qualitatively assess fidelity, we show Figure 914 (fMRI) and Figure 1518 (EEG) in Appendix O the generated samples closest to real ones in Frobenius distance. fMRI results are grouped by control (CN) and patient (non-CN); EEG by motor imagery class. DiffeoCFM reliably populates the neighborhood of real data, capturing class-conditional structure. We also show TriangCFM samples before projection, which appear realistic but are not SPD.
@@ 5.2 Neurophysiological Plausibility Study
Figure 2(a) shows group-level functional connectomes from the ADNI dataset, computed as the Fréchet mean (4) of correlation matrices conditioned on disease status. For comparison, Figure 8 in Appendix N presents the corresponding group-level connectomes derived from the ABIDE and OASIS3 datasets. The Fréchet mean is defined with respect to $\phi_{\text{Corr}_{d}}$ (12), the diffeomorphism used for the generation. For each class (CN and non-CN), we compare real connectomes (from held-out test subjects) to those generated by DiffeoCFM. In both cases, non-CN subjects exhibit reduced connectivity across hemispheres and between frontal and posterior regions—patterns commonly associated with mild cognitive impairment and Alzheimer’s disease [17].
@@
Figure 2(b) presents the group-level topographies of the first CSP filter across all subjects, derived from EEG generated by DiffeoCFM, alongside those from real EEG recordings in the BNCI2015-001 dataset. Subject-level topographies of the first CSP filter from the same dataset are further detailed in Figure 6 (Subjects 1–6) and Figure 7 (Subjects 7–12) in Appendix M, providing a detailed comparison across individuals. We visualize CSP spatial filters in two frequency bands associated with motor imagery. These filters highlight discriminative brain patterns distinguishing imagined right-hand versus feet movements.
@@
(b) Group-level topographic map using the first CSP’s spatial filter derived from real EEGs (BNCI2015-001) and generated data by DiffeoCFM Each map shows the first CSP’s spatial filter across $12$ subjects in the $\alpha$ ($8-12$ Hz) and $\beta$ ($13-30$ Hz) frequency bands during the first $2$ seconds following stimulus onset. Filters from DiffeoCFM at the group-level closely resemble real EEGs, preserving discriminative patterns of motor imagery classification.
@@
6 Conclusion
We introduced DiffeoCFM, an efficient framework for generating brain connectivity matrices that lie on Riemannian manifolds. By reformulating Riemannian flow matching through global diffeomorphisms, DiffeoCFM enabled fast training and sampling while ensuring manifold-constrained outputs by construction. Applied to both fMRI and EEG data, it produced realistic and neurophysiologically plausible samples, consistently outperforming existing baselines.
Nonetheless, several limitations remain. First, DiffeoCFM relies on a global diffeomorphism to Euclidean space, which exists for SPD and correlation matrices but not for all manifolds (e.g., Stiefel). Second, higher parcellation granularity leads to quadratic growth in dimension, increasing the amount of training data required. Third, common evaluation metrics like $\alpha$-precision and $\beta$-recall are geometry-agnostic and may miss neurophysiological structure.
Despite these challenges, generative modeling remains a promising direction for neuroimaging and BCI research. For instance, sharing trained generative models, rather than raw data, can facilitate multi-site collaboration with privacy guarantees [9]. To ensure both scientific reliability and practical impact, future work should develop geometry-aware evaluation metrics and integrate domain-specific validation protocols.
@@ Acknowledgment
This work was supported by grant ANR-22-PESN-0012 under the France 2030 program, managed by the Agence Nationale de la Recherche (ANR).
This work was granted access to the HPC resources of IDRIS under the allocation 2025-AD011016067 made by GENCI, and was carried out using the Jean Zay supercomputer.
Numerical computation was enabled by the scientific Python ecosystem: Matplotlib [31], Scikit-learn [50], Numpy [28], PyTorch [48], TorchCFM [58], PyRiemann [5], and MNE [26].
Appendix
@@ Appendix A Pullback Manifolds with Euclidean Spaces
This metric induces a Riemannian norm on the tangent space $T_{x}\mathcal{M}$ at each point $x\in\mathcal{M}$: $\|\xi\|_{x}=\sqrt{(\phi^{*}g_{E})_{x}(\xi,\xi)}$. The pair $(\mathcal{M},\phi^{*}g_{E})$ is then called a pullback manifold, and $\phi^{*}g_{E}$ is the pullback metric of $g_{E}$ by $\phi$. The pullback manifold is geodesically complete and admits globally unique geodesics [57, Chap. 7]. Moreover, many Riemannian operations—distances, geodesics, exponential and logarithmic maps, parallel transport, Fréchet means—reduce to simple computations in $E$. Given two points $x,y\in\mathcal{M}$, the geodesic $\gamma:[0,1]\to\mathcal{M}$ connecting $x$ to $y$ and its associated Riemannian distance are given by

### v2

@@ 4.3 Baselines
We present the baselines used for comparison with DiffeoCFM, including an oracle baseline (Real Data) that treats test data as if it were generated samples.
@@
These baselines apply generative models to the lower-triangular part of SPD or correlation matrices, a common heuristic in manifold modeling [52]. For fMRI, they use the strictly lower-triangular entries ($\phi:{\text{Corr}_{d}}\mapsto\mathbb{R}^{d(d-1)/2}$); for EEG, all lower-triangular entries ($\phi:{\mathbb{S}_{d}^{++}}\mapsto\mathbb{R}^{d(d+1)/2}$). TriangDDPM uses a DDPM [2], while TriangCFM trains a vector field using the standard CFM loss. Since $\phi$ is not a diffeomorphism, generated matrices that do not lie on the manifold are projected back onto it. In particular, generated matrices are not necessarily positive definite, so we apply a projection to ensure all eigenvalues are at least $\epsilon>0$ with $\mathbf{\Sigma}_{\text{proj}}=(1-\alpha)\mathbf{\Sigma}+\alpha\mathbf{I}$; see Appendix M for more details. These methods trade geometric fidelity for simplicity, relying on post hoc projections to enforce constraints.
RiemCFM
This method applies Riemannian CFM [12] directly on the SPD manifold using the affine-invariant metric [65]. The target conditional vector field is computed analytically from the Riemannian logarithm and exponential maps. Unlike TriangDDPM/TriangCFM, RiemCFM preserves the intrinsic geometry of $\mathbb{S}_{d}^{++}$} throughout training and sampling, yielding valid SPD matrices at every time step without post-hoc corrections. However, it requires computing geodesics, their derivatives and Riemannian norms under the affine-invariant metric, making it substantially more computationally expensive than all other presented methods. The reference implementation [^2] https://github.com/facebookresearch/riemannian-fm focuses exclusively on the SPD manifold under the affine-invariant metric, and does not provide a corresponding construction for correlation matrices.
@@
We apply the log–Euclidean map $\phi_{\mathbb{S}_{d}^{++}}$ (9) for EEG and the normalized Cholesky map $\phi_{\text{Corr}_{d}}$ (12) for fMRI in Algorithm 1 and 2. These diffeomorphisms allow Euclidean training with CFM while ensuring manifold-valid samples without post-processing.
Deep learning and training/sampling setups
TriangDDPM, TriangCFM and DiffeoCFM employ a two-layer MLP with $512$ hidden units, trained using AdamW [50] with a learning rate of $10^{-3}$ and batch size $64$. Training runs for $200$ epochs on fMRI and $2000$ epochs on EEG. RiemCFM has a 6-layer MLP with $512$ hidden units trained with AdamW (learning rate of $10^{-4}$), as recommended in [12]. These four methods use the dopri5 method from the torchdiffeq [11] library for time integration.
@@ 5 Results
We report both quantitative results and a neurophysiological plausibility study. Quantitative comparisons are summarized in Table 1, with additional analysis of projection effects in Table 2. Neurophysiological relevance is assessed in Figure 2, which shows class-conditional fMRI connectomes via Fréchet means and topographic maps of EEG Common Spatial Patterns (CSP) filters. To complement Table 1, we provide a visual summary of our findings in Figure 3 ( Appendix I). Together, these results show that DiffeoCFM produces realistic, class-conditional samples that preserve key features of brain connectivity. For a more detailed discussion of the baselines, please refer to Appendix H.
Table 1: Performance of generative models on 3 fMRI and 2 EEG datasets, evaluated with quality and Classification Accuracy Score (CAS) metrics. Quality metrics ($\alpha$-precision, $\beta$-recall, and $\alpha,\beta$-F1) assess alignment with the real distribution. CAS metrics (ROC-AUC and F1) evaluate downstream performance: a classifier is trained on generated data to predict disease status (fMRI) or motor imagery class (EEG), and tested on held-out real samples. Real Data rows use real samples to compare training and test distributions, serving as empirical upper bounds. The proposed method is denoted DiffeoCFM . $\text{mean}\,\pm\,\text{std}$ are reported. Bold values denote the best method and any methods that are not significantly worse than it (one-sided paired Wilcoxon signed-rank test, $\alpha=0.05$).
Quality Metrics CAS Metrics Time (s.) $\alpha$-Precision $\uparrow$ $\beta$-Recall $\uparrow$ $\alpha$$\beta$-F1 $\uparrow$ ROC-AUC $\uparrow$ F1 $\uparrow$ Training $\downarrow$ Sampling $\downarrow$ Dataset Method ABIDE Real Data 0.80  ± 0.08 0.79  ± 0.08 0.79  ± 0.03 0.67  ± 0.06 0.59  ± 0.07 N/A N/A DiffeoGauss 0.56  ± 0.06 0.29  ± 0.06 0.38  ± 0.06 0.66  ± 0.04 0.53  ± 0.06 0.07  ± 0.03 0.06  ± 0.00 TriangDDPM 0.04  ± 0.02 0.00  ± 0.00 0.00  ± 0.00 0.53  ± 0.06 0.47  ± 0.12 33.80  ± 1.19 0.37  ± 0.05 TriangCFM 0.04  ± 0.02 0.00  ± 0.00 0.00  ± 0.00 0.52  ± 0.05 0.40  ± 0.18 48.78  ± 1.27 0.79  ± 0.78 DiffeoCFM 0.77  ± 0.09 0.48  ± 0.07 0.59  ± 0.08 0.64  ± 0.06 0.58  ± 0.07 32.78  ± 0.96 0.40  ± 0.04 ADNI Real Data 0.91  ± 0.03 0.85  ± 0.06 0.88  ± 0.03 0.62  ± 0.05 0.62  ± 0.05 N/A N/A DiffeoGauss 0.02  ± 0.01 0.51  ± 0.08 0.04  ± 0.02 0.60  ± 0.05 0.29  ± 0.13 0.14  ± 0.01 0.18  ± 0.01 TriangDDPM 0.02  ± 0.00 0.00  ± 0.00 0.00  ± 0.00 0.53  ± 0.05 0.18  ± 0.11 90.03  ± 2.01 0.62  ± 0.08 TriangCFM 0.02  ± 0.00 0.00  ± 0.01 0.01  ± 0.01 0.56  ± 0.04 0.34  ± 0.13 87.37  ± 2.13 0.63  ± 0.09 DiffeoCFM 0.62  ± 0.11 0.77  ± 0.02 0.68  ± 0.06 0.63  ± 0.04 0.47  ± 0.10 88.01  ± 2.90 0.69  ± 0.11 OASIS-3 Real Data 0.88  ± 0.04 0.87  ± 0.03 0.88  ± 0.02 0.73  ± 0.05 0.63  ± 0.06 N/A N/A DiffeoGauss 0.51  ± 0.04 0.30  ± 0.04 0.38  ± 0.04 0.70  ± 0.05 0.41  ± 0.07 0.10  ± 0.01 0.13  ± 0.00 TriangDDPM 0.03  ± 0.01 0.00  ± 0.00 0.00  ± 0.00 0.54  ± 0.06 0.41  ± 0.14 70.39  ± 1.99 0.50  ± 0.06 TriangCFM 0.06  ± 0.01 0.00  ± 0.00 0.00  ± 0.00 0.52  ± 0.06 0.41  ± 0.14 67.92  ± 2.31 0.52  ± 0.07 DiffeoCFM 0.60  ± 0.05 0.35  ± 0.04 0.44  ± 0.04 0.67  ± 0.06 0.53  ± 0.07 67.83  ± 1.83 0.57  ± 0.05 BNCI 2014-002 Real Data 0.70  ± 0.05 0.60  ± 0.05 0.64  ± 0.03 0.83  ± 0.01 0.75  ± 0.02 N/A N/A DiffeoGauss 0.46  ± 0.04 0.77  ± 0.02 0.57  ± 0.03 0.80  ± 0.02 0.73  ± 0.02 0.06  ± 0.01 0.08  ± 0.01 TriangDDPM 0.43  ± 0.04 0.10  ± 0.02 0.16  ± 0.03 0.52  ± 0.03 0.20  ± 0.15 257.88  ± 0.14 0.30  ± 0.05 TriangCFM 0.48  ± 0.05 0.22  ± 0.03 0.30  ± 0.03 0.55  ± 0.03 0.24  ± 0.11 251.78  ± 0.85 0.35  ± 0.09 RiemCFM 0.67  ± 0.07 0.62  ± 0.06 0.63  ± 0.03 0.81  ± 0.02 0.72  ± 0.02 1983.58  ± 0.97 5.28  ± 0.58 DiffeoCFM 0.62  ± 0.04 0.63  ± 0.04 0.62  ± 0.02 0.81  ± 0.02 0.74  ± 0.02 253.04  ± 0.33 0.59  ± 0.08 BNCI 2015-001 Real Data 0.89  ± 0.01 0.89  ± 0.01 0.89  ± 0.00 0.73  ± 0.01 0.67  ± 0.01 N/A N/A DiffeoGauss 0.84  ± 0.01 0.90  ± 0.01 0.86  ± 0.01 0.73  ± 0.01 0.68  ± 0.01 0.07  ± 0.01 0.16  ± 0.01 TriangDDPM 0.73  ± 0.03 0.55  ± 0.03 0.63  ± 0.02 0.60  ± 0.02 0.59  ± 0.13 319.94  ± 5.88 0.37  ± 0.07 TriangCFM 0.79  ± 0.03 0.73  ± 0.03 0.76  ± 0.02 0.61  ± 0.02 0.59  ± 0.07 313.22  ± 2.16 0.38  ± 0.08 RiemCFM 0.93  ± 0.04 0.84  ± 0.05 0.88  ± 0.01 0.73  ± 0.01 0.66  ± 0.02 2753.93  ± 0.31 11.02  ± 0.59 DiffeoCFM 0.92  ± 0.01 0.86  ± 0.02 0.89  ± 0.01 0.73  ± 0.01 0.65  ± 0.01 319.83  ± 0.33 1.02  ± 0.08
@@ 5.1 Quantitative Study
[…]
Table 2: Impact of projection onto the manifolds $\mathbb{S}_{d}^{++}$} and $\text{Corr}_{d}$}: performance difference $\delta=$ TriangCFM - TriangCFM without projections. Negative $\Delta$ indicate degraded sample quality after projection onto the manifold. It enforces geometric constraints but severely reduces sample fidelity.
Dataset $\alpha$-precision $\beta$-recall $\alpha$$\beta$-F1 ABIDE -0.34 -0.69 -0.50 ADNI -0.63 -0.74 -0.69 OASIS-3 -0.52 -0.76 -0.64 BNCI2014-002 +0.13 -0.56 -0.19 BNCI2015-001 +0.00 -0.19 -0.09
@@
For the CAS metrics, which evaluate downstream predictive performance using ROC-AUC and F1 scores, DiffeoCFM consistently achieves strong results, often approaching the performance of Real Data. It obtains the highest ROC-AUC and F1 scores across all fMRI and EEG datasets. While DiffeoGauss remains competitive in terms of ROC-AUC, DiffeoCFM substantially outperforms it on F1 scores, with absolute gains of +0.05, +0.18, and +0.12 on ABIDE, ADNI, and OASIS-3, respectively. TriangDDPM and TriangCFM perform poorly across both CAS metrics on all datasets. This underperformance reflects the effect of the projection step, which alters generated matrices in ways detrimental to downstream classification. It is important to note that the classification pipeline requires inputs to be valid SPD matrices (i.e., elements of $\mathbb{S}_{d}^{++}$}). As a result, CAS metrics cannot be computed for TriangDDPM and TriangCFM without projection. For this reason, their performance without projection is not reported in Table 2.
@@
To qualitatively assess fidelity, we show Figure 1015 (fMRI) and Figure 1619 (EEG) in Appendix P the generated samples closest to real ones in Frobenius distance. fMRI results are grouped by control (CN) and patient (non-CN); EEG by motor imagery class. DiffeoCFM reliably populates the neighborhood of real data, capturing class-conditional structure. We also show TriangCFM samples before projection, which appear realistic but are not SPD.
@@ 5.2 Neurophysiological Plausibility Study
Figure 2(a) shows group-level functional connectomes from the ADNI dataset, computed as the Fréchet mean (4) of correlation matrices conditioned on disease status. For comparison, Figure 9 in Appendix O presents the corresponding group-level connectomes derived from the ABIDE and OASIS3 datasets. The Fréchet mean is defined with respect to $\phi_{\text{Corr}_{d}}$ (12), the diffeomorphism used for the generation. For each class (CN and non-CN), we compare real connectomes (from held-out test subjects) to those generated by DiffeoCFM. In both cases, non-CN subjects exhibit reduced connectivity across hemispheres and between frontal and posterior regions—patterns commonly associated with mild cognitive impairment and Alzheimer’s disease [19].
@@
Figure 2(b) presents the group-level topographies of the first CSP filter across all subjects, derived from EEG generated by DiffeoCFM, alongside those from real EEG recordings in the BNCI2015-001 dataset. Subject-level topographies of the first CSP filter from the same dataset are further detailed in Figure 7 (Subjects 1–6) and Figure 8 (Subjects 7–12) in Appendix N, providing a detailed comparison across individuals. We visualize CSP spatial filters in the $\alpha$ (8–12 Hz) and $\beta$ (13–30 Hz) bands that distinguish imagined right-hand from feet movements. The filters trained on real and on generated data concentrate on the same contralateral sensorimotor regions, mirroring the close CAS scores between Real data and DiffeoCFM in Table 1. This confirms that the generative model preserves the physiologically relevant information.
@@
(b) Group-level topographic map using the first CSP’s spatial filter derived from real EEGs (BNCI2015-001) and generated data by DiffeoCFM Each map shows the first CSP’s spatial filter across $12$ subjects in the $\alpha$ (8128-12 Hz) and $\beta$ (133013-30 Hz) frequency bands during the first 22 seconds following stimulus onset. Filters from DiffeoCFM at the group-level closely resemble real EEGs, preserving discriminative patterns of motor imagery classification.
@@
6 Conclusions, Limitations, and Future Works
We introduced DiffeoCFM, an efficient framework for generating brain connectivity matrices. By reformulating Riemannian flow matching through global diffeomorphisms, DiffeoCFM enabled fast training and sampling while ensuring manifold-constrained outputs by construction. Applied to fMRI and EEG data, it outperformed existing baselines with neurophysiologically plausible samples.
Nonetheless, several limitations remain. First, DiffeoCFM relies on a global diffeomorphism to Euclidean space, which exists for SPD and correlation matrices but not for all manifolds (e.g., Stiefel). Second, higher parcellation granularity makes the problem intrinsically hard: manifold dimension grows quadratically with region count, and the sample complexity grows exponentially [57]. Third, how alternative connectivity definitions (e.g., partial correlation or graphical-Lasso precision) affect generation quality remains an open question, since the choice of estimator directly defines the ground truth. Fourth, common evaluation metrics like $\alpha$-precision and $\beta$-recall are geometry-agnostic and may miss neurophysiological structure.
Despite these challenges, generative modeling remains a promising direction for neuroimaging and BCI research. For instance, sharing trained generative models, rather than raw data, can facilitate multi-site collaboration with privacy guarantees [10].
@@ Acknowledgment
This work was supported by grant ANR-22-PESN-0012 under the France 2030 program, managed by the Agence Nationale de la Recherche (ANR). CJ was supported by DATAIA Convergence Institute as part of the "Programme d’Investissement d’Avenir", (ANR-17-CONV-0003) operated by Inria. This work was performed using HPC resources from GENCI–IDRIS (Grant 2025-AD011016067).
Numerical computation was enabled by the scientific Python ecosystem: Matplotlib [33], Scikit-learn [61], Numpy [30], Scipy [72], PyTorch [59], fMRIprep [24], Nilearn [14], joblib [36], PyRiemann [5], torchcfm [69], torchdiffeq [11], pandas [58], and moabb [34].
@@ Appendix A Pullback Manifolds with Euclidean Spaces
This metric induces a Riemannian norm on the tangent space $T_{x}\mathcal{M}$ at each point $x\in\mathcal{M}$: $\|\xi\|_{x}=\sqrt{(\phi^{*}g_{E})_{x}(\xi,\xi)}$. The pair $(\mathcal{M},\phi^{*}g_{E})$ is then called a pullback manifold, and $\phi^{*}g_{E}$ is the pullback metric of $g_{E}$ by $\phi$. The pullback manifold is geodesically complete and admits globally unique geodesics [68, Chap. 7]. Moreover, many Riemannian operations—distances, geodesics, exponential and logarithmic maps, parallel transport, Fréchet means—reduce to simple computations in $E$. Given two points $x,y\in\mathcal{M}$, the geodesic $\gamma:[0,1]\to\mathcal{M}$ connecting $x$ to $y$ and its associated Riemannian distance are given by